Upon retirement, you hope to earn a monthly income of $41,984.36. The interest rate you will earn on your investment is 13.25%, and you expect to live for 22 years after retirement. How much must you have invested at the beginning of retirement so you achieve your target income and it lasts as long as you live?
Answer #1

Annuity payout formula

R = PV / ((1 - 1 / (1 + i)n) / i)
Where,
R = periodic payout
PV = Present Value
n = number of terms
i = period interest rate

Monthly income, R = $ 41984.36

Interest rate = 13.25% i.e 1.104% monthly

period = 22 years = 264 months

So PV = R * ((1 - 1 / (1 + i)n) / i)

=  41984.36 * ((1 - 1/(1.01104)264)/ 0.01104

= 3592934.96

investment required = $ 3,592,934.96